OpenStudy (anonymous):

Which expression is equivalent to sec^2x − 1? cot2x tan2x csc2x cos2x

OpenStudy (anonymous):

sec(x)=1/cos(x) so sec^2x=1/cos^2x now take the identity: sin^2x+cos^2x=1 divide all terms by cos^2x to get sin^2x/cos^2x+1=1/cos^2x then simplify using sin/cos = tan and 1/cos = sec tan^2x+1=sec^2x then just subtract 1 from each side sec^2x-1=tan^2x

OpenStudy (anonymous):

so b ?

OpenStudy (anonymous):

if b is \[\tan ^2(x) \] then yes